OpenStudy (anonymous):

PLS HELP WILL GIVE MEDAL AND FAN :) THX Question 1. 1. Factor. 16 - x2 (x - 4)(x + 4) (8 + x)(8 - x) (4 + x)(4 - x) (8 - x)(2 + x) Question 2. 2. Factor. 64a2 - 81. (32a -9)(32a + 9) (9 + 8a)(9 - 8a) (8a - 9)(8a - 9) (8a + 9)(8a - 9) Question 3. 3. Factor. 4x6 - 169. (2x3 + 13)(2x3 - 13) (2x5 + 13)(2x - 13) (2x + 13)(2x - 13) (2x3 + 13)(2x2 - 13) Question 4. 4. Factor. 9b8 -121.

OpenStudy (anonymous):

Question 4. 4. Factor. 9b8 121. (Points : 1) (3b4 11)(3b4 11) (3b4 + 11)(3b2 11) (3b4 + 11)(3b4 11) (3b2 + 11)(3b6 11) Question 5. 5. Factor. 81d4 1. (Points : 1) (9d2 + 1)(3d 1)(3d 1) (9d2 + 1)(3d + 1)(3d 1) (3d + 1)2(3d 1)2 (9d2 + 1)(3d + 1)

OpenStudy (sheraz12345):

Post ur questions one by one

OpenStudy (sheraz12345):

we are here to help u but not provide u the answers

OpenStudy (anonymous):

Anything that has a blank is a subraction sign

OpenStudy (sheraz12345):

for the first question u will use this formula

OpenStudy (anonymous):

16 - x2 would be factored into (4-x) and (4+x) To double-check, multiply the factors: (4-x)(4+x) = 16 + 4x - 4x -x2 = 16 - x2

OpenStudy (anonymous):

To factor 64a2 - 81, first step is to find a way to split 64a2 into two numbers. Next step is two split 81 into two numbers. Final step is to calculate the answer based on the first two steps. For example: 1. 8a * 8a = 64a2 2. 8a * -9 = -72a (outer) and 8a * 9 = 72 (inner) 3. -9 * 9 = -81 FINAL ANSWER : 64a2 - 72 + 72 - 81 = 64a2 - 81

OpenStudy (anonymous):

Wait so the awswer is what? lol

OpenStudy (anonymous):

sorry that was kinda confusing

OpenStudy (anonymous):

it ok

OpenStudy (anonymous):

i lost wht i was foing